# Prompt Guidance and Human Proximal Perception for HOT Prediction with Regional Joint Loss

### v1

@@ Abstract
The task of Human-Object conTact (HOT) detection involves identifying the specific areas of the human body that are touching objects. Nevertheless, current models are restricted to just one type of image, often leading to too much segmentation in areas with little interaction, and struggling to maintain category consistency within specific regions. To tackle this issue, a HOT framework, termed P3HOT, is proposed, which blends Prompt guidance and human Proximal Perception. To begin with, we utilize a semantic-driven prompt mechanism to direct the network’s attention towards the relevant regions based on the correlation between image and text. Then a human proximal perception mechanism is employed to dynamically perceive key depth range around the human, using learnable parameters to effectively eliminate regions where interactions are not expected. Calculating depth resolves the uncertainty of the overlap between humans and objects in a 2D perspective, providing a quasi-3D viewpoint. Moreover, a Regional Joint Loss (RJLoss) has been created as a new loss to inhibit abnormal categories in the same area. A new evaluation metric called “AD-Acc.” is introduced to address the shortcomings of existing methods in addressing negative samples. Comprehensive experimental results demonstrate that our approach achieves state-of-the-art performance in four metrics across two benchmark datasets. Specifically, our model achieves an improvement of 0.7$\uparrow$, 2.0$\uparrow$, 1.6$\uparrow$, and 11.0$\uparrow$ in SC-Acc., mIoU, wIoU, and AD-Acc. metrics, respectively, on the HOT-Annotated dataset. Code is available at https://github.com/YuxiaoWang-AI/P3HOT.
@@ 1 Introduction
Human-Object conTact (HOT) prediction [5] originated from Human-Object Interaction (HOI) detection, an emerging advanced semantic understanding task  [23, 25, 26]. HOI focuses on detecting humans, objects, and their interactions. It does not specify which parts of the human body come into contact with objects. However, identifying specific contact points on the human is required in the HOT task. This technology has a wide range of possible uses and can be applied in various fields such as human-computer interaction [23], virtual reality [12], and gesture recognition [6], opening up new opportunities in these areas.
HOT divides into 18 categories (including background), namely, Head, Chest, L_UpperArm, L_ForeArm, L_Hand, R_UpperArm, R_ForeArm, Buttocks, R_Hand, Hip, Back, L_Thigh, L_Calf, L_Foot, R_Thigh, R_Calf, R_Foot, and background, as shown in Figure 1(a). Chen et al. [5] introduced DHOT (using the encoder and decoder framework), which employed ResNet-50 as feature extraction network and designed a contact attention mechanism coupled with human masks for HOT prediction (Figure 1(b)). Nonetheless, DHOT only utilized features from the last layer of the feature extraction network, which can be disadvantageous for segmentation tasks that usually rely on low-level features to improve segmentation accuracy [6]. They predicted HOT from a single image modality, neglecting the guidance provided by other modalities. Additionally, the DHOT method failed to consider the consistency of classes within a region, leading to predictions that may include other classes within a particular class.
To overcome these challenges, we propose prompt guidance and human proximal perception method for HOT Prediction (a simplified diagram is shown in Figure 1(c)). Specifically, a text prompt mechanism is introduced to focus on human contact parts within the feature map. To filter out irrelevant information, a human proximal perception module is designed that utilizes a learnable parameter and human masks to dynamically perceive the depth range around the human. Furthermore, to regain initial texture detail loss during downsampling in the feature extraction stage, the lost image texture information is continuously refined by integrating outputs from each block of the feature extraction into the upsampling operations of the decoder. Finally, a novel loss function is designed, termed Regional Joint Loss (RJLoss), to ensure category consistency within regions and reduce the impact of abnormal categories.
Our contributions are summarized as follows:
@@
  • The proposed human proximal perception module is leveraged from a pseudo-3D perspective to assist HOT. A learnable threshold is used to dynamically perceive areas where interaction may exist, effectively reducing unnecessary interference.
  • A novel loss function, i.e., Regional Joint Loss, is introduced to ensure intra-region category consistency. A new evaluation metric is suggested to overcome the shortcomings of current metrics when accounting for incorrect predictions.
  • Extensive experiments have shown that our approach outperforms existing models on a newly introduced evaluation metric and achieves superior performance in SC-Acc., mIoU, and wIoU metrics on two benchmark datasets.
@@ 2 Related work
Figure 2: Overall architecture of the proposed P3HOT. The image encoder and text encoder are used to extract features from images and text, respectively. In the text encoder, we design a prompt template to create specific textual prompts. The similarity between the image and text features is computed to enhance the attention of the channel-specific feature maps in the image decoder. The human proximal perception module calculates the human mask and depth features for the input image, and uses learnable parameters, $\tau$, to capture the depth range around the human body and surrounding environment. This module serves as a filter on the feature maps. The image decoder progressively refines the segmentation features by integrating each block output from the image encoder, incorporating both textual and depth modality features, to predict the final segmentation map.
@@
HOT detection. Early HOT detection primarily focused on specific body parts such as hands and feet  [18, 13, 33, 6, 22, 31, 17]. DRNet [18] combines Mask R-CNN [11] to learn hand localization and predict its contact state. The foot contact detection is often used to estimate body posture and joint angles [13, 33]. These methods lack a holistic perspective, failing to provide a comprehensive assessment of HOT. To address this, Tripathi et al. [22] proposed a human contact detection method based on 3D scenes. However, this method suffers from low computational efficiency in 3D environments. Then two datasets are constructed based on 2D perspectives: HOT-Annotated and HOT-Generated. [5]. DHOT introduced a combination of ResNet and convolutional attention mechanisms to achieve HOT predictions [5]. Later, PIHOT [27] was designed to address the human-object occlusion problem. PIHOT employs a restoration model and a depth model to recover object features. However, these methods do not consider category consistency within a region, leading to frequent occurrences of other categories within a specific category. Additionally, PIHOT does not explore multimodal information to enhance contact detection performance. Multimodal information, such as text prompts, has been proven effective in various tasks  [15, 25, 26].
Why choose 2D over 3D contact detection? In many real-world applications, 2D tasks are sufficient for detecting HOT. For example, when understanding human-object interactions, it is sufficient to know whether the hands, feet, or other body parts are in contact with an object, without requiring precise 3D contact information. While 3D contact detection provides more information, it requires a larger number of parameters due to the complexity of 3D models, leading to slower speed. In contrast, 2D HOT detection is more suitable for some real-world applications. Moreover, 3D HOT detection datasets are currently scarce and challenging to construct. Although some recent works have attempted to create 3D HOT detection datasets [2, 22], the available data remains limited. These datasets are typically generated from existing 2D HOT datasets using the SMPL method [22], resulting in relatively coarse annotations for contact surfaces. Most importantly, 3D HOT detection datasets are only applicable to a limited range of objects [2], whereas 2D HOT datasets can be applied to a much broader variety of object categories [5].
@@ 3.2 Text Encoder
To extract text features, we first construct a text prompt template $T$, which takes the form: “A [body part] of the human body is in contact with an object.” Here, [body part] is sequentially replaced with each of the 17 human body parts (“Head”, “Chest”, “Left Upper Arm”, “Left Fore Arm”, “Left Hand”, “Right Upper Arm”, “Right Fore Arm”, “Buttocks”, “Right Hand”, “Hip”, “Back”, “Left Thigh”, “Left Calf”, “Left Foot”, “Right Thigh”, “Right Calf”, “Right Foot”) mentioned earlier. After encoding, we obtain $F_{TE}\in\mathbb{R}^{TN\times 1024}$, where $TN$ denotes the number of constructed text. The text encoder is initialized using the model proposed by Radford et al. [21], freezing its parameters during training. The similarity $\bm{S}\in\mathbb{R}^{1\times TN}$ between $\bm{F}_{IE}\in\mathbb{R}^{1\times 1024}$ and $\bm{F}_{TE}\in\mathbb{R}^{TN\times 1024}$ is then computed, via:
@@ 3.3 Human Proximal Perception
This module aims to preserve humans and the environment around them by eliminating irrelevant backgrounds. To accomplish this, the initial step is calculating the average depth for each person. Then, the depth range around each person is determined based on the learnable parameter $\tau$ to dynamically select the appropriate target region. Specifically, the improved SAM model is used to generate humans masks $\bm{M}$ based on the text prompt “person” [19], and apply the ZoeDepth model [1] to extract the depth features $\bm{D}\in\mathbb{R}^{H\times W}$ of the input image. Here, $\bm{M}=\{\bm{m}_{1},\bm{m}_{2},\dots,\bm{m}_{N}\}$, where $\bm{m}_{i}\in\mathbb{R}^{H\times W}$ represents the $i$-th human mask. In $\bm{m}_{i}$, 0 indicates the background, and 1 represents the human body. $N$ denotes the total number of persons. Next, we obtain the average depth of each human based on $\bm{M}$ and $\bm{D}$ as follows:
@@
$\bm{FM}=\begin{bmatrix}0&0&\cdots&0\\ 0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\end{bmatrix}_{H\times W},$ (8)
@@
Eq. 9 converts the depth values within the specified range to positive values, while those outside the range are set to negative values. The ReLU function in Eq. 10 keeps the original positive values and turns the negative values into zeros. Since the ReLU function is differentiable, it allows the parameter $\tau$ to be updated during backpropagation. DS denotes the downsampling operation, which reduces the size of $\bm{FM}$ from $H\times W$ to $\frac{H}{4}\times\frac{W}{4}$. $\oplus$ represents element-wise addition.
@@ 3.5 Loss
$\bm{O}_{c}(\text{ or }\bm{GT}_{c})=\{o_{pq}(\text{ or }gt_{pq})|p,q=1,2,\dots,\frac{H(W)}{4}\},$
$o_{pq}(\text{ or }gt_{pq})=\begin{cases}1,&\text{ if }o_{pq}(\text{ or }gt_{pq})=c\\ 0,&\text{ otherwise }\end{cases},$ (14)
@@
where CE and BE denote the cross-entropy loss and the binary cross-entropy loss, respectively. Both the image and text encoders are initialized with CLIP. We freeze the text encoder and train the image encoder using BE, enabling the image and text encoders to operate collaboratively. $\bm{S}\in\mathbb{R}^{1\times 18}$ represents the image-text matching similarity, and $\bm{\mathbb{C}}\in\mathbb{R}^{1\times 18}$ indicates the contact classes present in the input image. If class $c$ appears in the image, then $\bm{\mathbb{C}}_{c}=1$; otherwise, $\bm{\mathbb{C}}_{c}=0$.
@@ 4.2 Setup
For a fair comparison, ResNet-50 is used as the backbone of the image encoder module. In Eq. 21, $\alpha$, $\beta$, and $\gamma$ are set to 0.3, 0.1, and 1.0, respectively. The network is optimized using the AdamW optimizer. The batch size is set to 4 per GPU. The experimental environment is Ubuntu 20.04, equipped with 8 NVIDIA A6000 (48GB) GPUs. PyTorch version is 1.11.0, torchvision version is 0.12.0, and Python is 3.8.19.
@@ 4.3 Metrics
We found that the C-Acc. metric has some issues (see supplementary materials for details). If the prediction map classifies the entire image as a contact class, C-Acc. would be 100% because the predicted region encompasses the entire human body. This outcome is clearly undesirable, as most of the prediction pixels are incorrect in this case. To address this issue, we propose a new evaluation metric to replace C-Acc., called Adaptive Accuracy (AD-Acc.). We first extract the channel indices with the maximum values from the prediction map, resulting in a new $\bm{O}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}}$, where each element represents the predicted class. The calculation of AD-Acc. is then defined as follows:
$\bm{GT}^{B}(\text{ or }\bm{O}^{B})$ $=[gt_{ij}](\text{ or }[o_{ij}])$
$\bm{\zeta}=\bm{M}-\bm{GT}^{B}\otimes\bm{M},$ (23)
[…]
@@ 4.4 Performance
The experimental results (Table 1) demonstrate that our proposed model significantly outperforms all other models on both HOT-Annotated and HOT-Generated datasets across four evaluation metrics. For example, on the HOT-Annotated dataset, our model achieves a SC-Acc. of 46.0, which is notably higher than PIHOT at 45.3 and DHOT(ResNet-50)${}_{\text{Full}}$-OF at 40.1. Similarly, for mIoU, our model scores 25.6 compared to PIHOT at 23.6 and DHOT(ResNet-50)${}_{\text{Full}}$-OF at 22.1. Although our method is 5.8 points lower than PIHOT in terms of C-Acc., we previously mentioned that C-Acc. is not an accurate metric for evaluating the HOT task. This is because if every pixel in the entire image is predicted as a single contact category, C-Acc. would be 100, which is clearly not the desired outcome. To address this limitation, we propose a new evaluation metric, AD-Acc. In terms of AD-Acc., our method outperforms PIHOT by 11.0. On the HOT-Generated dataset, our model achieves a AD-Acc. of 30.6, surpassing PIHOT at 25.4 and DHOT(ResNet-50)${}_{\text{Full}}$-OF at 24.9. These substantial improvements emphasize the effectiveness of our approach and the critical role of fully integrated various mechanisms in achieving superior performance for HOT prediction tasks. Use more powerful encoder results, such as Swin-L [16]. See supplementary materials. In the supplementary materials, we include additional test results of the model, such as different encoders, inference time, and more.

### v2

@@ Abstract
The task of Human-Object conTact (HOT) detection involves identifying the specific areas of the human body that are touching objects. Nevertheless, current models are restricted to just one type of image, often leading to too much segmentation in areas with little interaction, and struggling to maintain category consistency within specific regions. To tackle this issue, a HOT framework, termed P3HOT, is proposed, which blends Prompt guidance and human Proximal Perception. To begin with, we utilize a semantic-driven prompt mechanism to direct the network’s attention towards the relevant regions based on the correlation between image and text. Then a human proximal perception mechanism is employed to dynamically perceive key depth range around the human, using learnable parameters to effectively eliminate regions where interactions are not expected. Calculating depth resolves the uncertainty of the overlap between humans and objects in a 2D perspective, providing a quasi-3D viewpoint. Moreover, a Regional Joint Loss (RJLoss) has been created as a new loss to inhibit abnormal categories in the same area. A new evaluation metric called “AD-Acc.” is introduced to address the shortcomings of existing methods in addressing negative samples. Comprehensive experimental results demonstrate that our approach achieves state-of-the-art performance in four metrics across two benchmark datasets. Specifically, our model achieves an improvement of 0.7$\uparrow$, 2.0$\uparrow$, 1.6$\uparrow$, and 11.0$\uparrow$ in SC-Acc., mIoU, wIoU, and AD-Acc. metrics, respectively, on the HOT-Annotated dataset. The sources code are available at https://github.com/YuxiaoWang-AI/P3HOT.
@@ 1 Introduction
Human-Object conTact (HOT) prediction [5] originated from Human-Object Interaction (HOI) detection, an emerging advanced semantic understanding task  [22, 24, 25]. HOI focuses on detecting humans, objects, and their interactions. It does not specify which parts of the human body come into contact with objects. However, identifying specific contact points on the human is required in the HOT task. This technology can be applied in various fields such as human-computer interaction [22], virtual reality [12], and gesture recognition [6].
HOT divides into 18 categories (including background) , as shown in Figure 1(a). Chen et al. [5] introduced DHOT, which employed ResNet-50 as feature extraction network and designed a contact attention mechanism coupled with human masks for HOT prediction (Figure 1(b)). Nonetheless, DHOT only utilized features from the last layer of the feature extraction network, which can be disadvantageous for segmentation tasks that usually rely on low-level features to improve segmentation accuracy [6]. They predicted HOT from a single image modality, neglecting the guidance provided by other modalities. Additionally, the DHOT method failed to consider the consistency of classes within a region, leading to predictions that may include other classes within a particular class.
To overcome these challenges, we propose prompt guidance and human proximal perception method for HOT Prediction (a simplified diagram is shown in Figure 1(c)). Specifically, a text prompt mechanism is introduced to focus on human contact parts within the feature map. A human proximal perception (HPP) module is designed that utilizes a learnable parameter and human masks to dynamically perceive the depth range around the human. Furthermore, to regain initial texture detail loss during downsampling in the feature extraction stage, the lost image texture information is continuously refined by integrating outputs from each block of the feature extraction into the upsampling operations of the decoder. Finally, a novel loss function is designed, termed RJLoss, to ensure category consistency within regions and reduce the impact of abnormal categories. Our contributions are summarized as follows:
@@
  • The proposed HPP module is leveraged from a pseudo-3D perspective to assist HOT. A learnable threshold is used to dynamically perceive areas where interaction may exist, effectively reducing unnecessary interference.
  • A novel loss function, i.e., RJLoss, is introduced to ensure intra-region category consistency. A new evaluation metric is suggested to overcome the shortcomings of current metrics when accounting for incorrect predictions.
@@ 2 Related work
Figure 2: Overall architecture. The image encoder and text encoder are used to extract features from images and text, respectively. In the text encoder, we design a prompt template to create specific textual prompts. The similarity between the image and text features is computed to enhance the attention in the image decoder. The HPP module calculates the human mask and depth features, and uses learnable parameters, $\tau$, to capture the depth range around the human body and surrounding environment. The image decoder progressively refines the segmentation features by integrating each block output from the image encoder.
@@
HOT detection. Early HOT detection primarily focused on specific body parts such as hands and feet  [17, 13, 32, 6, 21, 30, 16]. DRNet [17] combines Mask R-CNN [11] to learn hand localization and predict its contact state. The foot contact detection is often used to estimate body posture and joint angles [13, 32]. These methods lack a holistic perspective, failing to provide a comprehensive assessment of HOT. To address this, Tripathi et al. [21] proposed a human contact detection method based on 3D scenes. However, this method suffers from low computational efficiency in 3D environments. Then two datasets are constructed based on 2D perspectives: HOT-Annotated and HOT-Generated. [5]. DHOT introduced a combination of ResNet and convolutional attention mechanisms to achieve HOT predictions [5]. Later, PIHOT [26] was designed to address the human-object occlusion problem. PIHOT employs a restoration model and a depth model to recover object features. However, these methods do not consider category consistency within a region, leading to frequent occurrences of other categories within a specific category. Moreover, PIHOT has a large number of parameters, runs slowly, and does not exploit multimodal information to enhance contact detection performance.
Text-guided perception method. Multimodal information, such as text prompts, has been proven effective in various tasks  [15, 24, 25]. GEN-VLKT [15] leverages the CLIP model to extract text features, which are then used to initialize a fully connected layer, thereby improving accuracy. FreeA [24] employs the CLIP model to compute similarity between images and prompt texts, enabling the generation of candidate actions for label-free tasks. OpenCat [zheng2023open] enhances model performance through the use of text prompts. In contrast, our method utilizes text prompts to guide the network’s attention to 17 body parts (excluding the background), thereby improving segmentation accuracy by focusing more precisely on body parts.
Why choose 2D over 3D contact detection? In many real-world applications, 2D tasks are sufficient for detecting HOT. For example, when understanding human-object interactions, it is sufficient to know whether the hands, feet, or other body parts are in contact with an object, without requiring precise 3D contact information. While 3D contact detection provides more information, it requires a larger number of parameters due to the complexity of 3D models, leading to slower speed. In contrast, 2D HOT detection is more suitable for some real-world applications. Moreover, 3D HOT detection datasets are currently scarce and challenging to construct. Although some recent works have attempted to create 3D HOT detection datasets [2, 21], the available data remains limited. These datasets are typically generated from existing 2D HOT datasets using the SMPL method [21], resulting in relatively coarse annotations for contact surfaces. Most importantly, 3D HOT detection datasets are only applicable to a limited range of objects [2], whereas 2D HOT datasets can be applied to a much broader variety of object categories [5]. The number of datasets available for 2D HOT detection is gradually increasing, and their scope is also becoming broader.
@@ 3.2 Text Encoder
To extract texts features, we first construct a text prompts template $T$, which takes the form: “A [body part] of the human body is in contact with an object." Here, [body part] is sequentially replaced with each of the 17 human body parts (“Head", “Chest", “Left Upper Arm", “Left Fore Arm", “Left Hand", “Right Upper Arm", “Right Fore Arm", “Buttocks", “Right Hand", “Hip", “Back", “Left Thigh", “Left Calf", “Left Foot", “Right Thigh", “Right Calf", “Right Foot") mentioned earlier. After encoding, we obtain $\bm{F}_{TE}\in\mathbb{R}^{TN\times 1024}$, where $TN$ denotes the number of constructed texts. The text encoder is initialized using the model proposed by Radford et al. [20], freezing its parameters during training. The similarity $\bm{S}\in\mathbb{R}^{1\times TN}$ between $\bm{F}_{IE}\in\mathbb{R}^{1\times 1024}$ and $\bm{F}_{TE}\in\mathbb{R}^{TN\times 1024}$ is then computed, via:
@@ 3.3 Human Proximal Perception
This module aims to preserve humans and the environment around them by eliminating irrelevant backgrounds. To accomplish this, the initial step is calculating the average depth for each person. Then, the depth range around each person is determined based on the learnable parameter $\tau$ to dynamically select the appropriate target region. Specifically, the improved SAM model is used to generate humans masks $\bm{M}$ based on the text prompt “person" [18], and apply the ZoeDepth model [1] to extract the depth features $\bm{D}\in\mathbb{R}^{H\times W}$ of the input image. Here, $\bm{M}=\{\bm{m}_{1},\bm{m}_{2},\dots,\bm{m}_{N}\}$, where $\bm{m}_{i}\in\mathbb{R}^{H\times W}$ represents the $i$-th human mask. In $\bm{m}_{i}$, 0 indicates the background, and 1 represents the human body. $N$ denotes the total number of persons. Next, we obtain the average depth of each human based on $\bm{M}$ and $\bm{D}$ as follows:
@@
$\bm{FM}=[0]_{H\times W},$ (8)
@@
$[0]_{H\times W}$ denotes the initialization of a zero matrix of size $H\times W$. Eq. 9 converts the depth values within the specified range to positive values, while those outside the range are set to negative values. The ReLU function in Eq. 10 keeps the original positive values and turns the negative values into zeros. Since the ReLU function is differentiable, it allows the parameter $\tau$ to be updated during backpropagation. DS denotes the downsampling operation, which reduces the size of $\bm{FM}$ from $H\times W$ to $\frac{H}{4}\times\frac{W}{4}$. $\oplus$ represents element-wise addition. $\bm{O}$ denotes the fused feature map.
@@ 3.4 Image Decoder
The decoder network performs feature decoding. Specifically, it first upsamples the outputs from the four blocks of the encoder in sequence to fuse fine-grained features. Then, after integrating features from the HPP layer, it connects them through the attention output of the text encoder to predict the final segmentation map.
@@ 3.5 Loss
$\bm{O}_{c}(\bm{GT}_{c})=\{o_{pq}(gt_{pq})|p,q=1,2,\dots,\frac{H(W)}{4}\},$
$o_{pq}(gt_{pq})=\begin{cases}1,&\text{ if }o_{pq}(gt_{pq})=c\\ 0,&\text{ otherwise }\end{cases},$ (14)
@@
where CE and BE denote the cross-entropy loss and the binary cross-entropy loss, respectively. Both the image and text encoders are initialized with CLIP. We freeze the text encoder and train the image encoder using BE. $\bm{S}\in\mathbb{R}^{1\times 18}$ represents the image-text matching similarity, and $\bm{\mathbb{C}}\in\mathbb{R}^{1\times 18}$ indicates the contact classes present in the input image. If class $c$ appears in the image, then $\bm{\mathbb{C}}_{c}=1$; otherwise, $\bm{\mathbb{C}}_{c}=0$.
@@ 4.2 Setup
For a fair comparison, ResNet-50 is used as the backbone of the image encoder module. $\alpha$, $\beta$, and $\gamma$ are set to 0.3, 0.1, and 1.0, respectively. The network is optimized using the AdamW optimizer. The batch size is set to 4 per GPU. The experimental environment is Ubuntu 20.04, equipped with 8 NVIDIA A6000 GPUs. PyTorch version is 1.11.0, torchvision version is 0.12.0, and Python is 3.8.19.
@@ 4.3 Metrics
We found that the C-Acc. metric has some issues. If the prediction map classifies the entire image as a contact class, C-Acc. would be 100% because the predicted region encompasses the entire human body. This outcome is clearly undesirable, as most of the prediction pixels are incorrect in this case. To address this issue, we propose a new evaluation metric to replace C-Acc., called Adaptive Accuracy (AD-Acc.). We first extract the channel indices with the maximum values from the prediction map, resulting in a new $\bm{O}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}}$, where each element represents the predicted class. The calculation of AD-Acc. is then defined as follows:
$\bm{GT}^{B}(\text{ or }\bm{O}^{B})$ $=[gt_{ij}](\text{ or }[o_{ij}])$
$\bm{\zeta}=\bm{M}-\bm{GT}^{B}\otimes\bm{M},$ (23)
[…]
@@ 4.4 Performance
The experimental results (Table 1) demonstrate that our proposed model significantly outperforms all other models on both HOT-Annotated and HOT-Generated datasets across four evaluation metrics. For example, on the HOT-Annotated dataset, our model achieves a SC-Acc. of 46.0, which is notably higher than PIHOT at 45.3 and DHOT(ResNet-50)${}_{\text{Full}}$-OF at 40.1. Similarly, for mIoU, our model scores 25.6 compared to PIHOT at 23.6 and DHOT(ResNet-50)${}_{\text{Full}}$-OF at 22.1. Although our method is 5.8 points lower than PIHOT in terms of C-Acc., we previously mentioned that C-Acc. is not an accurate metric for evaluating the HOT task. This is because if every pixel in the entire image is predicted as a single contact category, C-Acc. would be 100, which is clearly not the desired outcome. To address this limitation, we propose a new evaluation metric, AD-Acc. In terms of AD-Acc., our method outperforms PIHOT by 11.0. On the HOT-Generated dataset, our model achieves a AD-Acc. of 30.6, surpassing PIHOT at 25.4 and DHOT(ResNet-50)${}_{\text{Full}}$-OF at 24.9. These substantial improvements emphasize the effectiveness of our approach and the critical role of fully integrated various mechanisms in achieving superior performance for HOT prediction tasks. Compared to PIHOI, our method is more computationally efficient, has fewer parameters, and lays the foundation for incorporating more modalities.